The perimeter of a rectangle is 34 ft. and the diagonal is 13 ft. What are the length and width of the rectangle?
Answer by mananth(16949)

2x+2y=34
2x=34-2y Divide equation by 2
x=17-y
x^2+y^2=13^2
(17-y)^2 +y^2=13^2

289 -34y+y^2+y^2= 169
2y^2-34y+120 =0
divide equation by 2
y^2-17y+60 = 0
y^2-12y-5y+60=0
y(y-12)-5(y-12)=0
(y-12)(y-5)=0
y=12 Or 5
The sides of the rectangle are 5 and 12